OpenStudy (anonymous):

Determine the type of intersection of the two planes 2x-5y+z-3=0 and 2x-y+5z-25=0?

OpenStudy (anonymous):

x = 61/4-(25/8)*z, y = 11/2-(5/4)*z, z = z solve for it

OpenStudy (anonymous):

thhere for it be a line